# Malle’s conjecture and Brauer groups of stacks

### v2

@@ 1.2. The Malle–Bhargava heuristics
Ellenberg and Venkatesh [40, Ques. 4.3] considered analogues of Malle’s conjecture where the discriminant is replaced by a different height function. This is quite natural even from the perspective of Malle’s conjecture, since for each transitive embedding $G\subseteq S_{n}$ one obtains a different counting problem. Moreover, numerous works have considered the problem of counting number fields with local conditions imposed. Bhargava [14] was the first to study this in the case of $S_{n}$-extensions of degree $n$. The expectation is that the leading constant should change in a simple way exactly corresponding in the local conditions imposed; in the literature this is often informally referred to as the Malle–Bhargava heuristics. For Bhargava’s original question see [15, §8.2], as well as Wood’s [93, §6.1] for a discussion of these heuristics and problems with the existing framework, and [92, §10] which notes that “An important open question is to even make a good conjecture about when exactly the principle should apply”.
@@ 1.3. Malle’s conjecture via stacks
    $\tau(k,G)$ $=\prod_{\begin{subarray}{c}v\in\operatorname{Val}(k)\\ v\mid\infty\end{subarray}}\frac{\#\operatorname{Hom}(\Gamma_{k_{v}},G)}{|G|}\prod_{\begin{subarray}{c}v\in\operatorname{Val}(k)\\ v\nmid\infty\end{subarray}}\frac{(1-1/q_{v})^{b(k,G)}}{|G|}\sum_{\varphi_{v}\in\operatorname{Hom}(\Gamma_{k_{v}},G)}\frac{1}{q_{v}^{v(\Delta_{\varphi_{v}})}}.$
@@
Borrowing terminology from Wood [91, §2.1], we call Case (1) the fair case. Here the factors which appear have precise analogues with Peyre’s constant [71, Def. 2.5]; for example $\tau(k,G)$ is a Tamagawa volume and $a(G)^{b(k,G)}/\#\widehat{G}(k)$ is a version of Peyre’s effective cone constant $\alpha$. However, the natural height to use in Manin’s conjecture for Fano varieties is the anticanonical height, and Peyre’s construction applies in this setting. Part of the reason for the pathological nature of $c_{\mathrm{Malle}}(k,G)$ is that the discriminant is not the anticanonical height in general; the role of the anticanonial height is played by what we call the radical discriminant $\operatorname{N}(\operatorname{rad}\Delta_{K/k})$, namely the norm of the ideal given by the product of ramified primes. As such Peyre’s formalism is insufficient in general, and we also require the framework of Batyrev and Tschinkel [11].
For Case (2), where the discriminant need not be fair, we instead consider the subgroup $M(G):=\langle\mathcal{M}(G)\rangle$ generated by the minimial index elements, which is normal as it is generated by conjugacy classes. We then sort homomorphisms according to the quotient map $G\to G/M(G)$; we make clear in our paper that this is a version of the Iitaka fibration from birational geometry, and this viewpoint is inspired by the work of Batyrev and Tschinkel [11]. This sorting corresponds to counting rational points in the fibres of the Iitaka fibration $BG\to B(G/M(G))$ then summing over all fibres. The first key observation is that the fibres of the Iitaka fibration can be identified with $BM(G)_{\varphi}$ where $M(G)_{\varphi}$ is a suitable inner twist of $M(G)$ (viewed as a group scheme). Thus one can interpret the count in the fibres as a version of Malle’s conjecture for group schemes. Our next key observation is that the restriction of the discriminant to a fibre now becomes a fair height function. This allows us to in fact obtain a precise prediction for the constants $c(k,G,\psi)$ which appear in Conjecture 1.3, with an analogous formula to the fair case. Altogether, this clarifies that the correct generality for Malle’s conjecture should allow $G$ to be a finite étale group scheme, so that one considers $G$-torsors instead of fields, and allows arbitrary height functions rather than just the discriminant. Our whole paper is written in this generality, including our most general conjectures in §9. This more general perspective to Malle’s conjecture is also taken in the papers [4, 5, 32, 33].
@@ 1.4.1. Mass formula, Hensel’s Lemma, and Tamagawa measures
In [14, Thm. 1.1] and [49, Prop. 5.3], Bhargava and Kedlaya prove formulae for weighted counts of number fields, which they call mass formulae. One of our first results concerns a generalisation of this to finite étale group schemes. To explain this we first introduce heights. We simplify slightly the exposition in the introduction, so some of the notation and set up appearing later in the paper is slightly different and more general (see §8.1 for the theory of heights).
@@
In §8.2 we define a local Tamagawa measure $\tau_{H,v}$ on $BG(k_{v})$ modelled on Peyre’s definition for Fano varieties [71, §2.2.1]. To obtain a global Tamagawa measure we define convergence factors $\lambda_{v}$ coming from the local Artin $L$-function factors of the collection $\mathcal{M}(w)=\{c\in\mathcal{C}_{G}:w(c)\neq 0\text{ is minimal}\}$ of minimal weight conjugacy classes of $G(-1)$. Our mass formula is used to prove that these indeed form convergence factors.
@@
With a measure in place, we are in a good position to put forward Conjecture 9.1 on the leading constant in Malle’s conjecture, as well as formalise the Malle–Bhargava heuristics through an equidistribution conjecture on the adelic points of $BG$. This is Conjecture 9.10, as well as the more general problem of imposing infinitely many local conditions, which leads to a property we call strong equidistribution (Conjecture 9.15). Crucially however in general one should restrict the measure to adelic points which is orthogonal to a suitable Brauer group; this extra factor does not appear in the Malle–Bhargava heuristics in the literature. It is necessary to take into account Grunwald–Wang-type phenomenon in the leading constant. We discuss the corresponding Brauer group in the next section.
@@
We give a definition of fair height functions, which generalises the special case of abelian groups by Wood [91] (Definition 8.1). This asks that the minimal weight conjugacy classes generate $G$. Our definition of fair is weaker than Wood’s in the case of abelian extensions. In fact Wood’s notion of fairness has strong implications for the corresponding Brauer groups which arise, so her choice can actually be explained through the Brauer group (see §10.7.4 for details). This framework allows us to answer Wood’s [92, §10] and Bhargava’s [15, §8.2] question on when the Malle–Bhargava heuristics should hold: this is exactly when counting by a fair height function and when there is no Brauer–Manin obstruction (providing one removes a possible accumulating collection of fields, as in Conjecture 1.3).
@@ 1.4.2. Brauer groups
Calculating this group $\operatorname{Br}_{\mathcal{C}}BG$ is crucial for calculating the leading constant in our conjecture. Firstly we prove that it is finite modulo $\operatorname{Br}k$ in the fair case, as required for the leading constant to even be well-defined; given that $\operatorname{Br}BG/\operatorname{Br}k$ is infinite in general, this finiteness is non-obvious. We also obtain an algorithm to calculate it as well as write down elements. We state this informally as a theorem here; see §6.7 for a precise description.
@@ 1.4.3. The exceptional set and the total count
The thinness of the breaking rational points in Manin’s conjecture [57, Thm. 1.4] is a difficult result proved used deep techniques in birational geometry. Our proof is substantially easier, and has the advantage of giving an explicit computable description of the breaking thin set, something which is currently lacking in the setting of Manin’s conjecture.
@@ 1.5. Structure of the paper
The next three sections form the technical heart of the paper. In §4 we introduce some of the key technical tools which we will use in our stacky framework, namely roots stacks, the cyclotomic inertia stack, and sectors. These are used to prove a new stacky version of Hensel’s lemma (Theorem 4.13). There already exist versions of Hensel’s lemma in the literature, which for a smooth stack $\mathcal{X}$ over a complete DVR $\mathcal{O}$ says that the map $\mathcal{X}(\mathcal{O})\to\mathcal{X}(\mathbb{F})$ to the residue field $\mathbb{F}$ is surjective, similarly to schemes (see e.g. [61, Lem. 4.2]). Our new version is completely different and has no analogue in the world of schemes. It gives a description of $\mathcal{X}(K)$ where $K$ is the fraction field of $\mathcal{O}$, and is based upon the modified valuative criterion for properness of stacks from [20].
@@ 1.6. Shortcut to the Conjecture:
The reader looking for a quick overview to understand the conjecture in the fair case should study §2.1 and §2.4 to get to grips with the definition of $BG$. One should then look at §3.1 to understand the Galois action on conjugacy classes and §3.4 for the definition of the Fujita invariant and the minimal weight conjugacy classes.
@@
One can now understand the conjectures in §9 regarding fair heights. A useful statement is Lemma 8.21, which shows that the leading constant in the conjecture is equal to an explicit finite sum of Euler products.
@@ 1.7. Notation and conventions
In the Malle’s conjecture literature, functions which assign a value to a number field go by various names, e.g. “generalised discriminant”, “counting function”, or “invariant”. We prefer the term height function to unify existing terminology regarding heights on varieties, and to avoid confusion with the local invariant $\operatorname{inv}_{v}:\operatorname{Br}k_{v}\to\mathbb{Q}/\mathbb{Z}$ from class field theory. A counting function for us is the function which gives the cardinality of the number of elements of bounded height.
@@ 2.1. The groupoid $BG(k)$
Let $G$ be an étale group scheme over a field $k$. The groupoid $BG(k)$ has multiple equivalent presentations. We follow the conventions for torsors as in [83, §2].
@@
(1) $\iff$ (3): The equivalence between this groupoid and $BG(k)$ follows from the fact that a cocycle is the same as a descent datum and the fact that descent for $G$-torsors is effective [84, Tag 0245].
@@ 2.2. Thin sets
Thin sets were introduced by Serre [81, §3.1] in his study of Hilbert’s irreducibility theorem. Peyre [72, §8] was the first to suggest that one should remove a thin set in Manin’s conjecture to avoid counter-example. This requirement is now part of all modern formulations of Manin’s conjecture. We will take a similar approach and remove a thin set in Malle’s conjecture to avoid counter-examples. Our definition of thin sets on $BG$ is as follows.
@@ 2.3. Inner twists
We now study inner twists and the automorphism groups of elements of $BG(k)$. Some of the results in this section are rephrasing in stacky terms properties of non-abelian Galois cohomology, as can be found in [80, §5].
@@
From a cohomological perspective, this is the twist of $G$ corresponding to the cohomology class $\varphi\in\operatorname{H}^{1}(k,G)$ where $G$ acts on itself by conjugation (see [80, Examples, §5.3] for more details).
@@
We next observe that inner twists do not change $BG$ (compare with [80, Prop. 35 bis]).
@@
For all $\varphi\in BG(k)$ we have $B\operatorname{\underline{Aut}}\varphi\cong BG$ [84, Tag 06QG]. However $\operatorname{\underline{Aut}}\varphi\cong G_{\varphi}$ by Lemma 2.9. ∎
@@
We next classify fibres of quotient maps with a rational point (compare with Corollary 2 in [80, §5.5]).
@@
Let $\varphi^{\prime}$ denote the image of $\varphi$ in the gerbe $f^{-1}(f(\varphi))$. Then $\operatorname{\underline{Aut}}\varphi^{\prime}=\ker(\operatorname{\underline{Aut}}\varphi\to\operatorname{\underline{Aut}}f(\varphi))=(\ker G_{\varphi}\to G_{\varphi}/H_{\varphi})=H_{\varphi}$. Thus by [84, Tag 06QG] we have $f^{-1}(f(\varphi))\cong B\operatorname{\underline{Aut}}\varphi^{\prime}\cong BH_{\varphi}$. ∎
@@
By [80, Cor. 5.5:2] the sum in the left-hand side is equivalent to a sum over the orbits of $G_{F(\varphi)}(k)$ acting on $G_{F(\varphi)}/H_{\varphi}(k)$. Moreover, by [80, Prop. 36], the stabiliser of this action for the orbit corresponding to $\varphi^{\prime}$ is $H_{\varphi^{\prime}}(k)$. Lemma 2.10 implies that $\operatorname{Aut}(\varphi^{\prime})=H_{\varphi^{\prime}}(k)$ and $\operatorname{Aut}(F(\varphi))=G_{F(\varphi)}(k)$. The desired statement then follows from the orbit-stabiliser theorem.
@@ 3.1. Galois action on conjugacy classes
The scheme $G(-1)$ arises in the following way. Heights will be determined by the restriction of a cocycle $\varphi:\Gamma_{k}\to G$ to the tame inertia group. The tame inertia is well-defined up to conjugacy, and is canonically isomorphic to a group of roots of unity [84, Tag 09EE]. Hence we canonically obtain an element of $G(-1)$, not of $G(\chi_{\mathrm{cycl}}^{-1})$ (see §7.1 and §8.1 for details). To help differentiate between $G(-1)$ and $G(\chi_{\mathrm{cycl}}^{-1})$, we will write $\gamma$ for an element of $G(-1)$ and $g$ for an element of $G(\chi_{\mathrm{cycl}}^{-1})$.
@@
For constant $G$ one can determine the Galois action on $G(\chi_{\mathrm{cycl}})(k^{\mathrm{sep}})/\mathrm{conj}$ through the character table of $G$, see [79, Ch. 12, Thm. 25]. In particular, the Galois action on $\mathcal{C}_{G}$ is trivial if and only if the character table of $G$ takes values in $k$.

### v3

@@ 1.2. The Malle–Bhargava heuristics
Ellenberg and Venkatesh [40, Ques. 4.3] considered analogues of Malle’s conjecture where the discriminant is replaced by a different height function. This is quite natural even from the perspective of Malle’s conjecture, since for each transitive embedding $G\subseteq S_{n}$ one obtains a different counting problem. Moreover, numerous works have considered the problem of counting number fields with local conditions imposed. Bhargava [14] was the first to study this in the case of $S_{n}$-extensions of degree $n$. The expectation is that the leading constant should change in a simple way exactly corresponding in the local conditions imposed; in the literature this is often informally referred to as the Malle–Bhargava heuristics. For Bhargava’s original question see [15, §8.2], as well as Wood’s [95, §6.1] for a discussion of these heuristics and problems with the existing framework, and [94, §10] which notes that “An important open question is to even make a good conjecture about when exactly the principle should apply”.
@@ 1.3. Malle’s conjecture via stacks
    $\tau(k,G)$ $=\prod_{\begin{subarray}{c}v\in\operatorname{Val}(k)\\ v\mid\infty\end{subarray}}\frac{\#\operatorname{Hom}(\Gamma_{k_{v}},G)}{|G|}\prod_{\begin{subarray}{c}v\in\operatorname{Val}(k)\\ v\nmid\infty\end{subarray}}\frac{(1-1/q_{v})^{b(k,G)}}{|G|}\sum_{\varphi_{v}\in\operatorname{Hom}(\Gamma_{k_{v}},G)}\frac{1}{q_{v}^{a(G)v(\Delta_{\varphi_{v}})}}.$
@@
In Case (1) we say that the discriminant is balanced. Here the factors which appear have precise analogues with Peyre’s constant [73, Def. 2.5]; for example $\tau(k,G)$ is a Tamagawa volume and $a(G)^{b(k,G)}/\#\widehat{G}(k)$ is a version of Peyre’s effective cone constant $\alpha$. However, the natural height to use in Manin’s conjecture for Fano varieties is the anticanonical height, and Peyre’s construction applies in this setting. Part of the reason for the pathological nature of $c_{\mathrm{Malle}}(k,G)$ is that the discriminant is not the anticanonical height in general; the role of the anticanonial height is played by what we call the radical discriminant $\operatorname{N}(\operatorname{rad}\Delta_{K/k})$, namely the norm of the ideal given by the product of ramified primes. As such Peyre’s formalism is insufficient in general, and we also require the framework of Batyrev and Tschinkel [11].
For Case (2), where the discriminant need not be balanced, we instead consider the subgroup $M(G):=\langle\mathcal{M}(G)\rangle$ generated by the minimial index elements, which is normal as it is generated by conjugacy classes. We then sort homomorphisms according to the quotient map $G\to G/M(G)$; we make clear in our paper that this is a version of the Iitaka fibration from birational geometry, and this viewpoint is inspired by the work of Batyrev and Tschinkel [11]. This sorting corresponds to counting rational points in the fibres of the Iitaka fibration $BG\to B(G/M(G))$ then summing over all fibres. The first key observation is that the fibres of the Iitaka fibration can be identified with $BM(G)_{\varphi}$ where $M(G)_{\varphi}$ is a suitable inner twist of $M(G)$ (viewed as a group scheme). Thus one can interpret the count in the fibres as a version of Malle’s conjecture for group schemes. Our next key observation is that the restriction of the discriminant to a fibre now becomes a balanced height function. This allows us to in fact obtain a precise prediction for the constants $c(k,G,\psi)$ which appear in Conjecture 1.3, with an analogous formula to the balanced case. Altogether, this clarifies that the correct generality for Malle’s conjecture should allow $G$ to be a finite étale group scheme, so that one considers $G$-torsors instead of fields, and allows arbitrary height functions rather than just the discriminant. Our whole paper is written in this generality, including our most general conjectures in §9. This more general perspective to Malle’s conjecture is also taken in the papers [4, 5, 32, 33].
@@ 1.4.1. Mass formula, Hensel’s Lemma, and Tamagawa measures
In [14, Thm. 1.1] and [49, Prop. 5.3], Bhargava and Kedlaya prove formulae for weighted counts of local fields, which they call mass formulae. One of our first results concerns a generalisation of this to finite étale group schemes. To explain this we first introduce heights. We simplify slightly the exposition in the introduction, so some of the notation and set up appearing later in the paper is slightly different and more general (see §8.1 for the theory of heights).
@@
In §8.2 we define a local Tamagawa measure $\tau_{H,v}$ on $BG(k_{v})$ modelled on Peyre’s definition for Fano varieties [73, §2.2.1]. To obtain a global Tamagawa measure we define convergence factors $\lambda_{v}$ coming from the local Artin $L$-function factors of the collection $\mathcal{M}(w)=\{c\in\mathcal{C}_{G}:w(c)\neq 0\text{ is minimal}\}$ of minimal weight conjugacy classes of $G(-1)$. Our mass formula is used to prove that these indeed form convergence factors.
@@
With a measure in place, we are in a good position to put forward Conjecture 9.1 on the leading constant in Malle’s conjecture, as well as formalise the Malle–Bhargava heuristics through an equidistribution conjecture on the adelic points of $BG$. This is Conjecture 9.10, as well as the more general problem of imposing infinitely many local conditions, which leads to a property we call strong equidistribution (Conjecture 9.15). Crucially however in general one should restrict the measure to those adelic points orthogonal to a suitable Brauer group; this extra factor does not appear in the Malle–Bhargava heuristics in the literature. It is necessary to take into account Grunwald–Wang-type phenomenon in the leading constant. We discuss the corresponding Brauer group in the next section.
@@
We give a definition of balanced height functions, which is modelled on the notion of fair counting function introduced by Wood for abelian $G$ [93, §2.1] (Definition 8.1). Balancedness asks that the minimal weight conjugacy classes generate $G$, whereas Wood asks that they generate the $r$-torsion of $G$ for all $r$. Wood’s notion of fairness has strong implications for the corresponding Brauer groups which arise, so her choice can actually be explained through the Brauer group (see §10.7.4 for details). This framework allows us to answer Wood’s [94, §10] and Bhargava’s [15, §8.2] question on when the Malle–Bhargava heuristics should hold: according to Conjecture 9.10 and §9.6, this should be exactly when counting by a balanced height function and when there is no Brauer–Manin obstruction (providing one removes a possible accumulating collection of fields, as in Conjecture 1.3).
@@ 1.4.2. Brauer groups
Calculating this group $\operatorname{Br}_{\mathcal{C}}BG$ is crucial for calculating the leading constant in our conjecture. Firstly we prove that it is finite modulo $\operatorname{Br}k$ in the balanced case, as required for the leading constant to even be well-defined; given that $\operatorname{Br}BG/\operatorname{Br}k$ is infinite in general, this finiteness is non-obvious. We also obtain an algorithm to calculate it as well as write down elements. We state this informally as a theorem here; see §6.7 for a precise description.
@@ 1.4.3. The exceptional set and the total count
The thinness of the breaking rational points in Manin’s conjecture [58, Thm. 1.4] is a difficult result proved used deep techniques in birational geometry. Our proof is substantially easier, and has the advantage of giving an explicit computable description of the breaking thin set, something which is currently lacking in the setting of Manin’s conjecture.
@@ 1.5. Structure of the paper
The next three sections form the technical heart of the paper. In §4 we introduce some of the key technical tools which we will use in our stacky framework, namely roots stacks, the cyclotomic inertia stack, and sectors. These are used to prove a new stacky version of Hensel’s lemma (Theorem 4.13). There already exist versions of Hensel’s lemma in the literature, which for a smooth stack $\mathcal{X}$ over a complete DVR $\mathcal{O}$ says that the map $\mathcal{X}(\mathcal{O})\to\mathcal{X}(\mathbb{F})$ to the residue field $\mathbb{F}$ is surjective, similarly to schemes (see e.g. [62, Lem. 4.2]). Our new version is completely different and has no analogue in the world of schemes. It gives a description of $\mathcal{X}(K)$ where $K$ is the fraction field of $\mathcal{O}$, and is based upon the modified valuative criterion for properness of stacks from [20].
@@ 1.6. Shortcut to the Conjecture:
The reader looking for a quick overview to understand the conjecture in the balanced case should study §2.1 and §2.4 to get to grips with the definition of $BG$. One should then look at §3.1 to understand the Galois action on conjugacy classes and §3.4 for the definition of the Fujita invariant and the minimal weight conjugacy classes.
@@
One can now understand the conjectures in §9 regarding balanced heights. A useful statement is Lemma 8.21, which shows that the leading constant in the conjecture is equal to an explicit finite sum of Euler products.
@@ 1.7. Notation and conventions
In the Malle’s conjecture literature, functions which assign a value to a number field go by various names, e.g. “generalised discriminant”, “counting function”, or “invariant”. We use the term height function to unify existing terminology regarding heights on varieties, and to avoid confusion with the local invariant $\operatorname{inv}_{v}:\operatorname{Br}k_{v}\to\mathbb{Q}/\mathbb{Z}$ from class field theory. A counting function for us is the function which gives the cardinality of the number of elements of bounded height.
@@ 2.1. The groupoid $BG(k)$
Let $G$ be an étale group scheme over a field $k$. The groupoid $BG(k)$ has multiple equivalent presentations. We follow the conventions for torsors as in [85, §2].
@@
(1) $\iff$ (3): The equivalence between this groupoid and $BG(k)$ follows from the fact that a cocycle is the same as a descent datum and the fact that descent for $G$-torsors is effective [86, Tag 0245].
@@ 2.2. Thin sets
Thin sets were introduced by Serre [83, §3.1] in his study of Hilbert’s irreducibility theorem. Peyre [74, §8] was the first to suggest that one should remove a thin set in Manin’s conjecture to avoid counter-examples. This requirement is now part of all modern formulations of Manin’s conjecture. We will take a similar approach and remove a thin set in Malle’s conjecture to avoid counter-examples. Our definition of thin sets on $BG$ is as follows.
@@ 2.3. Inner twists
We now study inner twists and the automorphism groups of elements of $BG(k)$. Some of the results in this section are rephrasing in stacky terms properties of non-abelian Galois cohomology, as can be found in [82, §5].
@@
From a cohomological perspective, this is the twist of $G$ corresponding to the cohomology class $\varphi\in\operatorname{H}^{1}(k,G)$ where $G$ acts on itself by conjugation (see [82, Examples, §5.3] for more details).
@@
We next observe that inner twists do not change $BG$ (compare with [82, Prop. 35 bis]).
@@
For all $\varphi\in BG(k)$ we have $B\operatorname{\underline{Aut}}\varphi\cong BG$ [86, Tag 06QG]. However $\operatorname{\underline{Aut}}\varphi\cong G_{\varphi}$ by Lemma 2.9. ∎
@@
We next classify fibres of quotient maps with a rational point (compare with Corollary 2 in [82, §5.5]).
@@
Let $\varphi^{\prime}$ denote the image of $\varphi$ in the gerbe $f^{-1}(f(\varphi))$. Then $\operatorname{\underline{Aut}}\varphi^{\prime}=\ker(\operatorname{\underline{Aut}}\varphi\to\operatorname{\underline{Aut}}f(\varphi))=(\ker G_{\varphi}\to G_{\varphi}/H_{\varphi})=H_{\varphi}$. Thus by [86, Tag 06QG] we have $f^{-1}(f(\varphi))\cong B\operatorname{\underline{Aut}}\varphi^{\prime}\cong BH_{\varphi}$. ∎
@@
By [82, Cor. 5.5:2] the sum in the left-hand side is equivalent to a sum over the orbits of $G_{F(\varphi)}(k)$ acting on $G_{F(\varphi)}/H_{\varphi}(k)$. Moreover, by [82, Prop. 36], the stabiliser of this action for the orbit corresponding to $\varphi^{\prime}$ is $H_{\varphi^{\prime}}(k)$. Lemma 2.10 implies that $\operatorname{Aut}(\varphi^{\prime})=H_{\varphi^{\prime}}(k)$ and $\operatorname{Aut}(F(\varphi))=G_{F(\varphi)}(k)$. The desired statement then follows from the orbit-stabiliser theorem.
@@ 3.1. Galois action on conjugacy classes
The scheme $G(-1)$ arises in the following way. Heights will be determined by the restriction of a cocycle $\varphi:\Gamma_{k}\to G$ to the tame inertia group. The tame inertia is well-defined up to conjugacy, and is canonically isomorphic to a group of roots of unity [86, Tag 09EE]. Hence we canonically obtain an element of $G(-1)$, not of $G(\chi_{\mathrm{cycl}}^{-1})$ (see §7.1 and §8.1 for details). To help differentiate between $G(-1)$ and $G(\chi_{\mathrm{cycl}}^{-1})$, we will write $\gamma$ for an element of $G(-1)$ and $g$ for an element of $G(\chi_{\mathrm{cycl}}^{-1})$.
@@
For constant $G$ one can determine the Galois action on $G(\chi_{\mathrm{cycl}})(k^{\mathrm{sep}})/\mathrm{conj}$ through the character table of $G$, see [81, Ch. 12, Thm. 25]. In particular, the Galois action on $\mathcal{C}_{G}$ is trivial if and only if the character table of $G$ takes values in $k$.